# Exploring Joint Observation of the CSST Shear and clustering of astrophysical gravitational wave source measurements

### v1

@@ 1 Introduction
The AGWB makes an important relation between galactic cosmology and compact object studies (G. Cusin et al., 2018), since the astrophysical gravitational wave sources (AGWS) can trace matter distribution through their clustering properties like galaxies (M. Oguri, 2016). Hence, the detection of AGWB or AGWS clustering will impose stringent constraints on compact object physics (e.g. equation of state and spin distributions), initial mass function, and cosmic star formation history (T. Regimbau, 2011), enable large-scale structure (LSS) investigations through anisotropy measurements, and provide novel constraints on cosmological parameters via angular power spectrum analysis (T. Namikawa et al., 2016; C. R. Contaldi, 2017; A. C. Jenkins et al., 2018; Y. Zheng et al., 2023).
On the other hand, according to Einstein’s general relativity, light propagating through inhomogeneous gravitational fields undergoes deflection. This weak gravitational lensing effect manifests as subtle distortions of distant source galaxies by foreground lensing structures (e.g. M. Bartelmann & P. Schneider, 2001). From these distorted images, we can reconstruct the intervening matter distribution along the line of sight, enabling precise mapping of the LSS (R. Mandelbaum, 2018).
Therefore, the cross-correlation between AGWS clustering and weak lensing signals can offer novel constraints on cosmological parameters, which would provide complementary insights to address current cosmological issues, such as the Hubble tension (A. G. Riess et al., 2021; Planck Collaboration et al., 2014; D. K. Hazra et al., 2015; B. Novosyadlyj et al., 2014; B. J. Luis et al., 2016) and $S_{8}$ tension (E. Di Valentino et al., 2021).
@@
The paper is organized as follows: in Section 2 we provide a detailed exposition of our theoretical framework of the AGWS clustering and weak lensing. In Section 3, we present the generation of the mock data from the three gravitational wave detector networks along with the CSST. The model fitting and constraint results are given in Section 4, and our conclusions are summarized in Section 5.
2 Methodology
2.1 AGWS clustering
@@
Here we adopt $\sigma_{{\rm ln}D}=0.05$ (S. Camera & A. Nishizawa, 2013). During their propagation from a source at redshift $z$ to Earth, gravitational waves are influenced by the LSS of the Universe, specifically through the weak gravitational lensing effect (R. Takahashi et al., 2005). Due to the magnification effect, the observed luminosity distance will be smaller than that predicted in a homogeneous and isotropic Friedmann-Lemaitre-Robertson-Walker (FLRW) cosmological model, which can be expressed as
@@
All gravitational wave sources are categorized into different distance or redshift bins based on the luminosity distance. The angular number density of gravitational wave sources $n_{i}^{gw}(\bm{\theta})$ in the $i$-th bin can be derived by projecting the 3D number density $n^{gw}(\bm{\theta},z)$ as (M. Oguri, 2016)
@@
The corresponding $S_{i}(z)$ and $T_{i}(z)$ are shown in Figure 1, when assuming the parameters of the ECS configuration (the details can be found in Section 3.1).
@@
where $T_{\rm obs}$ denotes the observational time of GW detectors, and $\dot{n}^{gw}(z)$ represents the detectable merger rate, calculated as the product of detector efficiency $E_{\rm det}(z)$ (the details can be found in Section 3.1) and intrinsic merger rate $R^{gw}(z)$, i.e. $\dot{n}^{gw}(z)=E_{\rm det}(z)\times R^{gw}(z)$. The $1+z$ factor accounts for cosmological time dilation. The first approximate equality in Equation (10) is due to the neglect of the second term in $G_{i}(z)$, as it is much smaller compared to the first term $S_{i}(z)$.
@@
We assume $\delta^{gw}(\bm{\theta},z)=b^{gw}(z)\delta_{\rm m}({\bm{\theta}},z)$, where $b^{gw}$ is the AGWS bias. We adopt a simple form $b^{gw}(z)=1+1/D(z)$ (M. Oguri, 2016), and D(z) is the linear growth factor. Then $\delta_{i}^{gw}(\bm{\theta})$ can be expressed as
@@ 2.2 Weak lensing
2.3 Angular Power Spectrum
$\times P_{\rm m}\left(k=\frac{\ell+1/2}{D_{\rm A}(z)},z\right),$ (26)
where $A,B=\{S,T,G,I\}$, and $P_{\rm m}(k,z)$ is the matter power spectrum calculated by CAMB (A. Lewis et al., 2000). Given gravitational wave localization uncertainties, we adopt $\ell_{\rm max}=300$ (K. Osato, 2018) and employ the linear matter power spectrum, neglecting the effects of neutrino and baryonic feedback.
$\hat{C}_{ij}^{gwwl}=(1+m_{j}){C}_{ij}^{gwwl}+N_{\rm sys}^{gwwl},$ (28)
3 Mock data
3.1 GW observation
@@
Figure 5: Detection rate comparison for the three detector networks. For BBH and BNS systems, we employ three-parameter sigmoid function fits (as listed in Table 1) to model their detection rates. For BHNS systems, we adopt the simplified assumption $E_{\rm det}^{\rm BHNS}(z)=1/2\left[E_{\rm det}^{\rm BBH}(z)+E_{\rm det}^{\rm BNS}(z)\right]$.
@@
Here $a$, $b$, and $c$ are the free parameters, which are given in Table 1 for the three detector networks. The ECS configuration shows superior performance in both detection counts and depth, followed by VKI+C and then HLKI+E. In Figure 5, we compare the detection rates for the three detector networks. We can find that, for the detection rate, we have ECS >> VKI+C >> HLKI+E, though HLKI+E outperforms VKI+C at low redshifts ($0<z<0.5$). Considering localization capabilities, we adopt $\ell_{\rm max}=300$ for ECS and $150$ for VKI+C and HLKI+E, when computing the cross-power spectra.
@@
Based on the merger rates of three compact binary types and detection efficiencies of the three detector networks, we select gravitational wave sources within the redshift range $[0,1.9]$, dividing them into 4 equal-number-count tomographic bins. For our baseline ECS configuration, each bin contains approximately $13,000$ compact binaries, while the VKI+C and HLKI+E networks have $5,400$ and $3,700$ sources per bin, respectively. We assume a constant bias $b_{i}^{gw}=1+1/D(z_{i}^{\rm eff})$ within each tomographic bin, where $z_{i}^{\rm eff}=\int_{z_{i}^{\rm min}}^{z_{i}^{\rm max}}z{n}^{gw}_{i}(z)dz$ denotes the effective redshift of the $i$-th bin, where ${n}^{gw}_{i}(z)=S_{i}(z)n^{gw}(z)$ is the GW sources distribution function for the $i$-th bin. $n^{gw}(z)=\frac{1}{\bar{n}_{i}^{gw}}\frac{cD_{\rm A}^{2}}{H(z)}\bar{n}^{gw}(z)$ is the total GW sources distribution function. We show the GW source total distribution function $n^{gw}(z)$ (black dotted curves) and distribution functions for the $i$-th bin $n^{gw}_{i}(z)$ (blue solid curves) in Figure 6, for the three GW dectection networks. It can be seen that the functional curve of the ECS configuration with the highest detection capability peaks at a higher redshift (left panel), and exhibits a greater peak value compared to the other two configurations (middle and right panels). The values of $z_{i}^{\rm eff}$, number $N_{i}^{gw}$, avarage angular number density $\bar{n}_{i}^{gw}$, and bias $b_{i}^{gw}$ of AGWS for the three detector networks in each bin are shown in Table 2.
3.2 CSST shear measurement
@@
The photometric redshift (photo-$z$) for individual galaxies is modeled with Gaussian probability distribution function (PDF) $z_{\rm obs}\sim N(z_{\rm true},\sigma_{z})$, where $\sigma_{z}$ is the true redshift and $\sigma_{z}=\sigma_{z0}(1+z)$ quantifies redshift uncertainty, assuming $\sigma_{z0}=0.05$ (Y. Gong et al., 2019). We divide the redshift range of the CSST photometric survey into four tomographic bins using equal-number intervals as listed in Table 3. Within the CSST survey area of $17,500\,\mathrm{deg}^{2}$, each bin contains approximately $4.1\times 10^{8}$ galaxies, with a total surface density of $26.7\;\mathrm{arcmin^{-2}}$. This result matches the prediction derived from the COSMOS catalog for the CSST photometric surveys (Y. Gong et al., 2019). We employ identical galaxy distributions for both lensing and source samples to better constrain the photo-$z$ uncertainty (see e.g. E. Schaan et al., 2020).
Figure 8: The covariance matrix for the joint data vector of the AGWS clustering, AGWS-weak lensing and cosmic shear.
@@
For the normalized redshift distributions $n_{i}^{g}$ of lens and source galaxies in each tomographic bin, we introduce the shift parameter $\Delta z_{i}$ and stretch parameter $\sigma_{i}^{z}$ to account for potential uncertainties (R. Cawthon et al., 2022; A. Porredon et al., 2022). Then the galaxy redshift distribution for each tomographic bin is given by
@@
3.3 Covariance matrix and Mock Data
@@
$\mathrm{Cov}_{ijmn}^{ABCD}$ $=\mathrm{Cov}[\hat{C}_{ij}^{AB}(\ell),\;\hat{C}_{mn}^{CD}(\ell^{\prime})]=\frac{\delta_{\ell\ell^{\prime}}}{(2\ell+1)f_{\rm sky}\Delta\ell}$
@@
where $f_{\rm sky}$ is the sky coverage fraction, $A,B,C,D\in\{gw,wl\}$ denote different tracers, and $i,j,m,n\in\{1,2,3,4\}$ indicate tomographic bins. We adopt $N_{\ell}=15$ logarithmic multipole bins within $\ell\in[10,300]$, and set $\ell_{\rm max}=300$ for ECS network, $\ell_{\rm max}=150$ for VKI+C and HLKI+E networks and $\ell_{\rm min}=10$ to ensure the validity of Limber’s approximation. The computed covariance matrix is shown in Figure 8.
@@
4 Constraint and Result
4.1 Fitting Method
@@
Table 4: The fiducial values and flat priors of the 30 free parameters considered in our constraint process.
@@
AGWS bias (for ECS)
$b_{1}^{gw}$ 2.400 $\mathcal{U}(1.8,3.8)$
$b_{2}^{gw}$ 2.695 $\mathcal{U}(1.8,3.8)$
$b_{3}^{gw}$ 2.924 $\mathcal{U}(1.8,3.8)$
$b_{4}^{gw}$ 3.170 $\mathcal{U}(1.8,3.8)$

### v2

@@ I Introduction
The astrophysical gravitational wave sources (AGWS) makes an important relation between galactic cosmology and compact object studies (G. Cusin et al., 2018), since the AGWS can trace matter distribution through their clustering properties like galaxies (M. Oguri, 2016). Hence, the detection of AGWS clustering will impose stringent constraints on compact object physics (e.g. equation of state and spin distributions), initial mass function, and cosmic star formation history (T. Regimbau, 2011), enable large-scale structure (LSS) investigations through anisotropy measurements, and provide novel constraints on cosmological parameters via angular power spectrum analysis (T. Namikawa et al., 2016; C. R. Contaldi, 2017; A. C. Jenkins et al., 2018; Y. Zheng et al., 2023).
The cross-correlation between AGWS with LSS tracers such as galaxies represents a transformative multi-messenger probe (A. Balaudo et al., 2023; S. Mukherjee et al., 2024), which can effectively extract cosmological information (S. Mukherjee et al., 2021; G. Sala et al., 2025). In addition, this approach enables the measurement of the AGWS clustering bias (S. Zazzera et al., 2025), with the tomographic analysis of the cross-correlation signal across redshift bins, which could reveal the evolution of the bias and offer a direct test of structure formation and merger rate evolution (F. Calore et al., 2020). The inclusion of relativistic effects such as magnification further allows tests of gravity on cosmological scales and provides independent constraints on the AGWS population properties (S. Mukherjee et al., 2020; G. Scelfo et al., 2023; S. Afroz & S. Mukherjee, 2024).
On the other hand, the weak gravitational lensing survey provides a direct probe of the projected matter distribution and its growth over cosmic time. In modern cosmological analyses, this signal is quantified via the cosmic shear angular power spectrum (or its real-space counterpart, the two-point correlation function), which is sensitive primarily to the parameters of amplitude of matter fluctuations $\sigma_{8}$ and the matter density $\Omega_{m}$ (e.g. M. Bartelmann & P. Schneider, 2001). The Stage-III and Stage-IV surveys always employ a multi-probe $3\times 2\rm{pt}$ framework (E. Krause et al., 2021; A. Dvornik et al., 2023; D. Sciotti et al., 2024) that jointly analyzes three two-point functions of weak lensing and galaxy clustering, which can extract cosmological information accurately and break the degeneracies between parameters, such as DES Y6 (DES Collaboration et al., 2026) and DESI DR1 $3\times 2\rm{pt}$ analyses (A. Porredon et al., 2025).
Compared to the AGWS–galaxy cross-correlation, the cross-correlation between AGWS and weak lensing can directly link the spatial distribution of GW events to the underlying matter field. Such a measurement is sensitive not only to the expansion history but also to the growth of structure. Consequently, the cross-correlation between AGWS and weak lensing can deliver novel and self-consistent cosmological constraints, which would provide complementary insights to address current cosmological issues, such as the Hubble tension(A. G. Riess et al., 2021; Planck Collaboration et al., 2014; D. K. Hazra et al., 2015; B. Novosyadlyj et al., 2014; B. J. Luis et al., 2016), and explore the properties of dark matter, dark energy, and AGWS population.
@@
The paper is organized as follows: in Section II we provide a detailed exposition of our theoretical framework of the AGWS clustering and weak lensing. In Section III, we present the generation of the mock data from the three gravitational wave detector networks along with the CSST. The model fitting and constraint results are given in Section IV, and our conclusions are summarized in Section V.
II Methodology
II.1 AGWS clustering
@@
Here we adopt $\sigma_{{\rm ln}D}=0.05$ (S. Camera & A. Nishizawa, 2013). During their propagation from a source at redshift $z$ to Earth, gravitational waves are influenced by the LSS of the Universe, specifically through the weak gravitational lensing effect (R. Takahashi et al., 2005). Due to the weak gravitational lensing effect, the observed luminosity distance to a source is statistically modified relative to the value predicted in a homogeneous and isotropic Friedmann-Lemaitre-Robertson-Walker (FLRW) cosmological model. This modification arises from the integrated matter distribution along the line of sight, which can be expressed as
@@
All gravitational wave sources can be categorized into different bins based on the luminosity distance, and their redshifts can be derived by assuming or giving a specific cosmology. The angular number density of gravitational wave sources $n_{i}^{gw}(\bm{\theta})$ in the $i$-th bin can be derived by projecting the 3D number density $n^{gw}(\bm{\theta},z)$ as (M. Oguri, 2016)
@@
The corresponding $S_{i}(z)$ and $T_{i}(z)$ are shown in Figure 1, when assuming the parameters of the ECS configuration (the details can be found in Section III.1).
@@
where $T_{\rm obs}$ denotes the observational time of GW detectors, and $\dot{n}^{gw}(z)$ represents the detectable merger rate, calculated as the product of detector efficiency $E_{\rm det}(z)$ (the details can be found in Section III.1) and intrinsic merger rate $R^{gw}(z)$, i.e. $\dot{n}^{gw}(z)=E_{\rm det}(z)\times R^{gw}(z)$. The $1+z$ factor accounts for cosmological time dilation. The first approximate equality in Equation (10) is due to the neglect of the second term in $G_{i}(z)$, as it is much smaller compared to the first term $S_{i}(z)$.
@@
We assume $\delta^{gw}(\bm{\theta},z)=b^{gw}(z)\delta_{\rm m}({\bm{\theta}},z)$, where $b^{gw}$ is the AGWS bias. We adopt a simple form $b^{gw}(z)=1+1/D(z)$ (M. Oguri, 2016) , and D(z) is the linear growth factor. Note that, when generating the mock data, we treat the bias in each tomographic bin as a constant calculated using this formula (see Section III.1 for details), and set it as a free parameter in the fitting process. Then $\delta_{i}^{gw}(\bm{\theta})$ can be expressed as
@@ II.2 Weak lensing
II.3 Angular Power Spectrum
$\times P_{\rm m}\left(k=\frac{\ell+1/2}{D_{\rm A}(z)},z\right),$ (26)
where $A,B=\{S,T,G,I\}$, and $P_{\rm m}(k,z)$ is the matter power spectrum calculated by CAMB (A. Lewis et al., 2000). Given gravitational wave localization uncertainties, we adopt $\ell_{\rm max}=300$ (K. Osato, 2018) (only for ECS, the corresponding $k_{\rm max}$ is between 0.09 and 0.18 ${\rm Mpc^{-1}}h$ at the effective redshifts of the tomographic bins we consider, see Section III.1 for details), and employ the linear matter power spectrum, neglecting the effects of neutrino and baryonic feedback.
$\hat{C}_{ij}^{gwwl}=(1+m_{j}){C}_{ij}^{gwwl}+N_{\rm sys}^{gwwl},$ (28)
III Mock data
III.1 GW observation
@@
Figure 5: Detection rate comparison for the three detector networks. For BBH and BNS systems, we employ three-parameter sigmoid function fits (as listed in Table 1) to model their detection rates.
@@
Here $a$, $b$, and $c$ are the free parameters, which are given in Table 1 for the three detector networks. The ECS configuration shows superior performance in both detection counts and depth, followed by VKI+C and then HLKI+E.
In Figure 5, we compare the detection rates for the three detector networks. Note that, for the BHNS systems, we adopt a simplified assumption $E_{\rm det}^{\rm BHNS}(z)=1/2\left[E_{\rm det}^{\rm BBH}(z)+E_{\rm det}^{\rm BNS}(z)\right]$, due to large astrophysical uncertainties in the BHNS formation rates and the lack of standardized forecasts for the third-generation detectors. Our results are insensitive to this assumption, since the BHNS systems are expected to constitute a small fraction of the total AGWS catalog, and their contribution to the clustering signal is subdominant compared to the BNS systems. We can find that, for the detection rate, we have ECS >> VKI+C >> HLKI+E, though HLKI+E outperforms VKI+C at low redshifts ($0<z<0.5$). Considering localization capabilities, we adopt $\ell_{\rm max}=300$ for ECS and $150$ for VKI+C and HLKI+E, when computing the cross-power spectra.
@@
Based on the merger rates of three compact binary types and detection efficiencies of the three detector networks, we select gravitational wave sources within the redshift range $[0,1.9]$, dividing them into 4 equal-number-count tomographic bins. For our baseline ECS configuration, each bin contains approximately $13,000$ compact binaries, while the VKI+C and HLKI+E networks have $5,400$ and $3,700$ sources per bin, respectively. When generating the mock data, we assume a constant bias $b_{i}^{gw}=1+1/D(z_{i}^{\rm eff})$ within each tomographic bin, where $z_{i}^{\rm eff}=\int_{z_{i}^{\rm min}}^{z_{i}^{\rm max}}z{n}^{gw}_{i}(z)dz$ denotes the effective redshift of the $i$-th bin, where ${n}^{gw}_{i}(z)=S_{i}(z)n^{gw}(z)$ is the GW sources distribution function for the $i$-th bin. $n^{gw}(z)=\frac{1}{\bar{n}_{i}^{gw}}\frac{cD_{\rm A}^{2}}{H(z)}\bar{n}^{gw}(z)$ is the total GW sources distribution function. We show the GW source total distribution function $n^{gw}(z)$ (black dotted curves) and distribution functions for the $i$-th bin $n^{gw}_{i}(z)$ (blue solid curves) in Figure 6, for the three GW dectection networks. It can be seen that the functional curve of the ECS configuration with the highest detection capability peaks at a higher redshift (left panel), and exhibits a greater peak value compared to the other two configurations (middle and right panels). The values of $z_{i}^{\rm eff}$, number $N_{i}^{gw}$, avarage angular number density $\bar{n}_{i}^{gw}$, and bias $b_{i}^{gw}$ of AGWS for the three detector networks in each bin are shown in Table 2.
III.2 CSST shear measurement
@@
The photometric redshift (photo-$z$) for individual galaxies is modeled with Gaussian probability distribution function (PDF) $z_{\rm obs}\sim N(z_{\rm true},\sigma_{z})$, where $\sigma_{z}$ is the true redshift and $\sigma_{z}=\sigma_{z0}(1+z)$ quantifies redshift uncertainty, assuming $\sigma_{z0}=0.05$ (Y. Gong et al., 2019). While this parametrization captures the dominant contribution of photo-$z$ uncertainty, it neglects higher-order features commonly present in real photo-$z$ probability distribution, such as non-Gaussian tails, catastrophic outliers, and biases induced by template mismatch or calibration errors. These may propagate into the tomographic window functions and bias the inferred angular power spectra. In weak lensing analyses, it typically manifests as a smoothing of the cross-bin correlations and a slight degradation in cosmological parameter constraints (Z. Ma et al., 2006; D. Huterer et al., 2006; J. Myles et al., 2021). However, in our forecast analysis, this simplification is reasonable and sufficient, and it would not affect our mian conclusions.
We divide the redshift range of the CSST photometric survey into four tomographic bins using equal-number intervals as listed in Table 3. Within the CSST survey area of $17,500\,\mathrm{deg}^{2}$, each bin contains approximately $4.1\times 10^{8}$ galaxies, with a total surface density of $26.7\;\mathrm{arcmin^{-2}}$. This result matches the prediction derived from the COSMOS catalog for the CSST photometric surveys (Y. Gong et al., 2019). We employ identical galaxy distributions for both lensing and source samples to better constrain the photo-$z$ uncertainty (see e.g. E. Schaan et al., 2020).
Figure 8: As an example, the covariance matrix at $\ell\simeq 11$ for the joint data vector of the AGWS clustering, AGWS-weak lensing and cosmic shear power spectra for different tomographic bin combinations.
@@
For the normalized redshift distributions $n_{i}^{g}$ of lens and source galaxies in each tomographic bin, we introduce the shift parameter $\Delta z_{i}$ and stretch parameter $\sigma_{i}^{z}$ to account for potential uncertainties (R. Cawthon et al., 2022). The shift parameter $\Delta z_{i}$ accounts for an additive shift to the mean redshift of each bin due to the global bias in the photo-$z$ calibration, and the stretch parameter $\sigma_{i}^{z}$ denotes the uncertainty on the width of the redshift distribution (A. Porredon et al., 2022). Then the galaxy redshift distribution for each tomographic bin is given by
@@
III.3 Covariance matrix and Mock Data
@@
$\mathrm{Cov}_{ijmn}^{ABCD}(\ell)$ $=\mathrm{Cov}[\hat{C}_{ij}^{AB}(\ell),\;\hat{C}_{mn}^{CD}(\ell^{\prime})]=\frac{\delta_{\ell\ell^{\prime}}}{(2\ell+1)f_{\rm sky}\Delta\ell}$
@@
where $f_{\rm sky}$ is the sky coverage fraction, $A,B,C,D\in\{gw,wl\}$ denote different tracers, and $i,j,m,n\in\{1,2,3,4\}$ indicate tomographic bins. We adopt $N_{\ell}=15$ logarithmic multipole bins within $\ell\in[\ell_{\rm min},\ell_{\rm max}]$, and set $\ell_{\rm max}=300$ for ECS network and $\ell_{\rm max}=150$ for VKI+C and HLKI+E networks. This choice reflects the different angular resolution and source localization capabilities of each network. Networks with longer baselines and more geographically distributed sites achieve better sky localization, enabling reliable reconstruction of small-scale (high $\ell$) clustering modes. We set $\ell_{\rm min}=10$ to ensure the validity of Limber’s approximation. As an example, the computed covariance matrix at $\ell\simeq 11$ is shown in Figure 8.
@@
IV Constraint and Result
IV.1 Fitting Method
@@
Table 4: The fiducial values and flat priors of the 30 free parameters considered in our constraint process for ECS/VLI+C/HLKI+E.
@@
AGWS bias (for ECS/VKI+C/HLKI+E)
$b_{1}^{gw}$ 2.400/2.359/2.318 $\mathcal{U}(1.8,3.8)$
$b_{2}^{gw}$ 2.695/2.647/2.599 $\mathcal{U}(1.8,3.8)$
$b_{3}^{gw}$ 2.924/2.886/2.853 $\mathcal{U}(1.8,3.8)$
$b_{4}^{gw}$ 3.170/3.152/3.142 $\mathcal{U}(1.8,3.8)$